40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 10 00:09:33 2012, kiwamu, Update, LSC, spike hunting in REFL33 
    Reply  Tue Jan 17 01:02:07 2012, kiwamu, Update, LSC, glitch hunting in REFL RFPDs : strange PRMsingle_bounce.png
       Reply  Tue Jan 17 02:27:49 2012, kiwamu, Update, LSC, fringe tests : all the suspensions are innocent 
       Reply  Tue Jan 17 13:47:40 2012, kiwamu, Update, LSC, dirty steering mirror in the REFL path 
          Reply  Wed Jan 18 12:36:26 2012, kiwamu, Update, LSC, wiped a steering mirror on the REFL path DSC_3861_small.jpgDSC_3863_small.jpg
       Reply  Tue Jan 17 19:07:47 2012, rana, Update, LSC, glitch hunting in REFL RFPDs : strange 
       Reply  Wed Jan 18 14:28:36 2012, kiwamu, Update, LSC, estimation of optical length between PRM and scattering object calibration.pngscattering.png
          Reply  Wed Jan 18 16:31:10 2012, kiwamu, Update, LSC, estimation of optical length between PRM and scattering object ND1atten.png
Message ID: 6183     Entry time: Tue Jan 10 00:09:33 2012     Reply to this: 6202
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: spike hunting in REFL33 

[John / Kiwamu]

 We tried to figure out what is causing spikes in the REFL33 signal, which is used to lock PRCL.

No useful information was obtained tonight and it is still under investigation.

 


(Background)

 One thing preventing us from doing smooth measurements of the noise budget and the sensing matrix is some sharp spikes in the LSC error signals.

For example when we lock PRMI with REFL33 and AS55 fedback to PRCL and MICH respectively, both the REFL33 and AS55 signals show some spikes in time series.

Those spikes then bring the noise spectra higher than how they should be.

So for the reason, taking the noise budget doesn't give us much information about the interferometer rather than there are spikes.

Also the sensing matrix measurement has been suffered from those spikes, which excite the impulse responses of the low pass filters in the LOCKIN detection systems a lot.

 

(What we did)

 We looked into the actual analog signals to see if there are indeed spikes or not before they are acquired to the ADCs.

But we didn't find any corresponding spikes in the signals that are after the mixers.

It maybe because the signals we looked into didn't have high enough SNR because they were coming out from the monitor lemo outputs on the demod boards.

 Then we thought the spikes are from the whitening circuits, due to some kind of saturation.

We decreased the gain of the whitening filters by a factor of 10, but it didn't help and the spikes were still there.

ELOG V3.1.3-